Amira knows that the time it takes her to commute to work is approximately normally distributed with a mean of 45 minutes and a standard deviation of 3 minutes. What time must she leave home in the morning so that she is 95% sure of arriving at work by 9.00am? 



1
Expert's answer
2021-02-02T04:44:55-0500

Assume that random variable "X" denotes a time that Amira needs to arive at work.

We need to find such a minimum number "\alpha" that "P(X\leq\alpha)=0.95". We remind that the density fo the normal distribution with parameters "\mu=45" and "\sigma=3" is: "p(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12(\frac{x-\mu}{\sigma})^2}=\frac{1}{3\sqrt{2\pi}}e^{-\frac12(\frac{x-45}{3})^2}" . Thus, "P(X\leq\alpha)=\int_{-\infty}^{\alpha}\frac{1}{3\sqrt{2\pi}}e^{-\frac12(\frac{x-45}{3})^2}dx". The latter yields "0.95" in case "\alpha=49.94". We used the following code in Anaconda to get the result:

from scipy import integrate

import numpy as np

import math


func = lambda x:(1/(3*math.sqrt(2)*math.sqrt(math.pi)))*math.exp(-1/2*((x-45)/3)*((x-45)/3))


e = integrate.quad(func, -np.inf, 49.94)

print(e)


We subsituted different values in the command "integrate.quad" to get the result

Thus, it is enough to leave at 8.10 (more precisely, at 8.10 and 6 seconds) to arrive at work till 9.00am.